OpenStudy (cutegirl1000):

For which operations is the set {–1, 0, 1} closed? Choose all answers that are correct. A. addition B. division C. multiplication D. subtraction

OpenStudy (cutegirl1000):

Can anyone please HELP!!!!!!|

zepdrix (zepdrix):

So we look at `any two elements in the set`, and perform the given operation, If we `end up with` an element outside of our set, then we DO NOT have closure.

OpenStudy (cutegirl1000):

I don't get this topic at all! I tried really hard and I couldn't do it... this is 8th grade math and I'm in 6th!!!

zepdrix (zepdrix):

Let's look at addition first. -1 + 0 = -1 I took the elements -1 and 0 from the set, performing addition, we ended up with -1, which is in the set, so it looks like we might have closure, ya? We need to check every combination of numbers though. We're really just looking to see if there is any problems, -1 + 0 = -1 -1 + 1 = 0 -1 + -1 = -2 1 + 1 = 2 0 + 1 = 1

zepdrix (zepdrix):

Notice that when using addition, we were able to come up with -2 and 2. These are values outside of our set. So we see that we DO NOT have closure.

OpenStudy (cutegirl1000):

I think the answers may be multiplication & division

zepdrix (zepdrix):

Mmmm close, division actually gives us a problem though. Recall that you can't divide by 0 in the land of math :) 1/0 does not exist. So this operation is not well defined for our set.